# Bridging Privacy and Robustness for Trustworthy Machine Learning

### v4

@@ IV-A2 PAC Robustness Measurement
Consequently, $\mathcal{M}$ is termed $\alpha$-PAC-Robust with probability $1-\beta$.
This definition essentially evaluates the consistency of the algorithm’s performance against variations in the input data. A robust algorithm remains accurate even when small portions of the dataset are altered, thus reflecting resilience against data perturbations. This definition motivates the design of an algorithm $\mathcal{A}(\mathcal{D})$ that is sufficiently accurate (and potentially randomized) for estimating a parameter $\mu(P)$, given that the input dataset $\mathcal{D}$ is drawn from a specific distribution $P$.
Both the privacy and robustness measurements presented here play instrumental roles in developing and assessing algorithms within the intricately balanced domain of data utility and confidentiality. The nuanced understanding of these metrics enables practitioners to formulate strategies that concurrently enhance the security and integrity of the analytical processes while mitigating risks to data privacy.
IV-B Theoretical Connections between Privacy Notions
In the following, relations among MBP, ABP and LDP are investigated. Through mathematical definitions and theorems, we demonstrate how these concepts of privacy are interconnected and how one can transition from one privacy framework to another. Specifically, we will prove through theorems that an algorithm satisfying the requirements of LDP also satisfies those of MBP, and vice versa. Additionally, we will discuss the connection between MBP and ABP, along with how to convert from a mapping that provides maximum privacy protection to one that offers an average level of privacy protection. These theoretical results not only deepen our understanding of the connections between various privacy measures but also offer significant guidance for crafting and analyzing algorithms that preserve privacy.
LDP aims to safeguard individual data contributors’ privacy, enabling the extraction of aggregate statistics or insights from the data. The concept of LDP is articulated as follows.
Definition IV.7 ($\epsilon$-LDP).
A randomized function $\mathcal{M}$ defined as $\epsilon$-local differentially private if, $\forall$ possible datasets $d$ and $d^{\prime}$, $forall$ subset S of the output space, the following inequality is true
$\frac{\Pr[\mathcal{M}(d)\in S]}{\Pr[\mathcal{M}(d^{\prime})\in S]}\leq e^{\epsilon}.$ (5)
In this formula, $\epsilon$ is non-negative, which quantifies the level of privacy protection, and $\Pr[\mathcal{M}(d)\in S]$ represents the probability that the randomized mechanism $\mathcal{M}$ outputs a result in the set S when given the dataset $d$.
This definition states that the probability of any outcome being produced by the mechanism when given dataset $d$ should be similar to the probability of the same outcome being produced when given a dataset $d^{\prime}$, smaller $\epsilon$ values indicate stronger privacy guarantees. LDP ensures that individual contributions to the dataset are protected by adding noise or randomness to the results of queries or computations, making it difficult to infer specific details about any individual in the dataset.
The following Lemma IV.1 implicates that if an algorithm provides $\xi$-LDP, then it also offers $\xi$-MBP. In other words, the same mechanism that provides local privacy protection, where individual data points are protected, also ensures a certain level of global privacy in a Bayesian sense. This means that the mechanism is not only useful for protecting individual data points but also for safeguarding the privacy of the dataset as a whole. Please check C-A for detailed analysis and proof.
Lemma IV.1.
Let MBP be defined using Definition IV.3. Consider a case where the adversary’s prior belief distribution follows a uniform distribution. That is, $f_{D}(d)=f_{D}(d^{\prime})$, $\forall d^{\prime},d\in\mathcal{D}$. If $f_{W|D}(\cdot|\cdot)$ is $\xi$-LDP. That is, the mapping $f_{W|D}(\cdot|\cdot)$ satisfies that $\forall d^{\prime},d\in\mathcal{D}$,
$\frac{f_{W|D}(w|d)}{f_{W|D}(w|d^{\prime})}\in[e^{-\xi},e^{\xi}],$
then $f_{W|D}(\cdot|\cdot)$ constitutes a privacy-preserving transformation ensuring $\xi$-Maximum Bayesian Privacy, where for all $d\in\mathcal{D}\ and\ w\in\mathcal{W}$,
$\left|\log\left(\frac{f_{D|W}(d|w)}{f_{D}(d)}\right)\right|\leq\xi.$ (6)
Conversely, the following lemma states that if an algorithm offers $\xi$-MBP, it also provides $2\xi$-LDP. This suggests that mechanisms crafted to safeguard the privacy of the dataset as a whole, as evaluated by Bayesian privacy, also afford a degree of protection for individual data points within a local differential privacy framework.
Lemma IV.2.
Let MBP be defined using Definition IV.3. Consider a situation where the adversary’s prior belief distribution follows a uniform distribution. Consequently, $f_{D}(d)=f_{D}(d^{\prime})$ for any $d,d^{\prime}$. If an algorithm is $\xi$-Maximum Bayesian Privacy, then it is $(2\xi)$-differentially private. Suppose privacy-preserving mapping $f_{W|D}(\cdot|\cdot)$ ensures $\xi$-maximum Bayesian privacy. That is, $\xi=\max\limits_{w\in\mathcal{W},d\in\mathcal{D}}\left|\log\left(\frac{f_{D|W}(d|w)}{f_{D}(d)}\right)\right|$. Then, the function $f_{W|D}(\cdot|\cdot)$ achieves $(2\xi)$-differential privacy
$\frac{f_{W|D}(w|d)}{f_{W|D}(w|d^{\prime})}\in[e^{-2\xi},e^{2\xi}].$
This lemma establishes a bridge between local differential privacy and maximum Bayesian privacy, showing that mechanisms satisfying one privacy criterion also offer a certain level of privacy protection according to the other criterion. This provides insights into the relationship between different privacy definitions and can be useful when designing and analyzing privacy-preserving algorithms in various contexts.
With Lemma IV.1 and Lemma IV.2, we are now ready to derive the relationship between MBP and LDP.
Theorem IV.3 (Theoretical Relationship between MBP and LDP).
Consider a situation in which the adversary’s prior belief distribution follows a uniform distribution. Let MBP be defined using Definition IV.3. If an algorithm is $\xi$-LDP, then it is $\xi$-MBP. If an algorithm is $\xi$-MBP, then it is $2\xi$-LDP.
The first conclusion of Theorem IV.4 addresses a scenario in which the prior belief distribution of an adversary closely resembles a uniform distribution, meaning that the adversary has a relatively unbiased and uniform belief about the data points in the set $d$. The lemma introduces an algorithm with a privacy-preserving mapping denoted as $f_{W|D}(\cdot|\cdot)$. This mapping provides a maximum Bayesian privacy guarantee, where privacy is quantified by the parameter $\xi$. The parameter $\xi$ quantifies the maximum privacy loss resulting from the mapping, determined by the logarithm of the ratio between the conditional posterior distribution of a data point $d$ given an observation $w$ and the prior distribution of $d$. In other words, $\xi$ measures the maximum privacy leakage under this mechanism.
The second conclusion of Theorem IV.4 implies that if the algorithm provides $\xi$-maximum Bayesian privacy, then it also provides $(2\xi)$-differential privacy. This implies that the privacy-preserving mapping not only demonstrates robustness in terms of Bayesian privacy but also fulfills a corresponding concept of privacy, namely, differential privacy. The differential privacy guarantee of mapping is expressed in terms of the ratios of conditional probabilities. Specifically, the ratio of the conditional probabilities of observing $w$ given two different data points $d$ and $d^{\prime}$ falls within the range of $[e^{-2\xi},e^{2\xi}]$. This range specifies the extent to which the mapping distorts the likelihood of observing $w$ given different data points. The larger $\xi$ is, the wider this range becomes, indicating stronger privacy protection. This conclusion shows a connection between maximum Bayesian privacy and differential privacy. It illustrates that a mapping ensuring maximum Bayesian privacy, particularly when the adversary’s prior belief distribution approximates uniformity, also confers a degree of differential privacy. This finding holds significance in comprehending and quantifying the privacy assurances offered by such mappings within the realm of privacy-preserving data analysis.
We could also explore the situation where the prior belief distribution of the adversary is $\epsilon$-close to a uniform distribution and the analytical approach would be akin to the current scenario.
Theorem IV.4 (Generalized Relationship between MBP and LDP).
Consider the scenario where the prior belief distribution of the adversary is "$\epsilon$-close" to a uniform distribution, namely $\forall d\text{ and }d^{\prime},\ |\log\frac{f_{D}(d)}{f_{D}(d^{\prime})}|\leq\epsilon$. Let MBP be defined using Definition IV.3. If an algorithm is $\xi$-LDP, then it is $(\xi+\epsilon)$-MBP. If an algorithm is $\xi$-MBP, then it is $(2\xi+\epsilon)$-LDP.
Remark: We might also explore the situation where the prior belief distribution of the adversary is $\epsilon$-close to a uniform distribution and the analytical approach would be akin to the current scenario.
The following theorem demonstrates the relationship between average Bayesian privacy and maximum Bayesian privacy.
Theorem IV.5 (Relationship Between MBP and ABP).
@@ IV-B Theoretical Connections between Privacy Notions
Let MBP be defined using Definition IV.3. For any $\epsilon_{p,m}\geq 0$, suppose privacy-preserving mapping $f_{W|D}(\cdot|\cdot)$ ensuring $\epsilon_{p,m}$-Maximum Bayesian Privacy. Namely,
@@
where $\epsilon_{p,a}$ is introduced in Definition IV.1.
The following theorem demonstrates the generalized relationship between average Bayesian privacy and maximum Bayesian privacy. This theorem establishes a generalized relationship between MBP and ABP under the assumption that the ratio of the attacker’s prior belief $f^{\mathcal{B}}_{D}(d)$ and the true prior distribution $f_{D}(d)$ is bounded within the range $[e^{-\epsilon},e^{\epsilon}]$. The theorem states that if a privacy-preserving mapping $f_{W|D}(\cdot|\cdot)$ guarantees $\epsilon_{p,m}$-MBP, i.e., the ratio of the posterior belief $f_{D|W}(d|w)$ and the prior belief $f_{D}(d)$ is bounded within the range $[e^{-\epsilon_{p,m}},e^{\epsilon_{p,m}}]$ for any $w\in\mathcal{W}$ and $d\in\mathcal{D}$, then the ABP (denoted as $\epsilon_{p,a}$) is bounded by a function of $\epsilon_{p,m}$ and $\epsilon$. Specifically, the theorem provides an upper bound on $\epsilon_{p,a}$ in terms of $\epsilon_{p,m}$ and $\epsilon$, which can be used to analyze the relationship between MBP and ABP under certain conditions on the attacker’s prior belief. This result is useful for understanding the connections between different notions of Bayesian privacy and can be leveraged to design privacy-preserving algorithms with desired properties.
Theorem IV.6 (Generalized Relationship Between MBP and ABP).
Assume that the prior belief of the attacker satisfies that
$\frac{f^{\mathcal{B}}_{D}(d)}{f_{D}(d)}\in[e^{-\epsilon},e^{\epsilon}].$
Let MBP be defined using Definition IV.3. For any $\epsilon_{p,m}\geq 0$, let $f_{W|D}(\cdot|\cdot)$ be a privacy preserving mapping that guarantees $\epsilon_{p,m}$-maximum Bayesian privacy. That is,
$\frac{f_{D|W}(d|w)}{f_{D}(d)}\in[e^{-\epsilon_{p,m}},e^{\epsilon_{p,m}}],$
where $\epsilon_{p,a}$ is introduced in Definition IV.1.
IV-C PAC Robustness of Private Algorithms
This subsection establishes a significant result linking the $(\epsilon,\beta,\alpha)$-MBP privacy criterion to PAC robustness. It proves that any algorithm satisfying MBP inherently possesses a quantifiable level of robustness against data perturbations, with probabilistic guarantees. This finding bridges the gap between theoretical privacy and practical robustness, offering insights for designing machine learning models that are both privacy-preserving and resilient.
The following Theorem IV.7 states that an algorithm $\mathcal{M}$, conforming to the $(\epsilon,\beta,\alpha)$-MBP privacy criterion, exhibits a specified degree of robustness against perturbations in the dataset. The robustness metric is defined as $(\frac{\log(1/\gamma)}{2n\epsilon},\alpha)$, representing the bounds within which the algorithm’s output is expected to vary. This level of robustness is probabilistically assured with a confidence of $1-\beta/\gamma$. The parameter $\gamma$ functions as a tunable confidence level within the exclusive range of 0 and 1, while $n$ quantifies the dataset’s size, and $\beta$ indicates the algorithm’s permissible error rate.
Theorem IV.7 (PAC Robustness of $\epsilon$-MBP Protection Mechanisms).
Let $\mathcal{M}$ be an $(\epsilon,\beta,\alpha)$-MBP (Definition IV.5) algorithm. Then $\mathcal{M}$ is $\alpha$-PAC-Robust (Definition IV.6) with probability $1-\gamma$, where $\gamma=(1+4\epsilon)\beta$, $n$ represents the size of the dataset, and $\beta$ represents the error probability.
This theorem elucidates a significant implication for privacy-preserving algorithms in machine learning, specifically those adhering to the $(\epsilon,\beta,\alpha)$-MBP model. It articulates that any algorithm $\mathcal{M}$ satisfying this MBP criterion inherently possesses a quantifiable level of PAC robustness. This robustness is defined in terms of $\alpha$ and $\gamma$, where $\gamma$ is a parameter that influences the confidence level, $n$ is the dataset size, and $\beta$ is the permitted error probability. Crucially, the theorem assures that with probability $1-\beta/\gamma$, the algorithm will maintain its robust performance despite potential variations in the data, thereby offering a probabilistic guarantee of robustness in the face of data perturbations. This bridges the gap between theoretical privacy guarantees and practical robustness in applied machine learning models.
This theorem is significant as it establishes a formal connection between robustness in the presence of learning errors and measurable criteria for privacy protection. It sets the stage for exploring the crucial relationship between data privacy and algorithmic robustness in machine learning systems. It highlights the importance of understanding and quantifying this connection to develop algorithms that simultaneously protect sensitive information and maintain performance in the face of perturbations or adversarial attacks. The theorem motivates the need for theoretical frameworks and metrics to analyze the privacy-robustness trade-off.
V Conclusions
In this paper, we have explored the intricate relationships between different privacy notions, namely LDP, ABP, and MBP, and their implications for algorithmic robustness. We established theoretical connections between these privacy concepts, proving that LDP implies MBP and vice versa under certain conditions, and deriving bounds between MBP and ABP. These findings provide valuable insights for constructing privacy-preserving algorithms. We also investigated the relationship between PAC robust learning and privacy preservation, demonstrating how to derive PAC robustness from privacy-preserving algorithms and construct privacy-preserving algorithms from PAC robust ones. This connection highlights the interplay between privacy and robustness in machine learning. Our findings contribute to the growing body of research on privacy and robustness in machine learning, providing a deeper understanding of the connections between different privacy notions and their implications for algorithmic robustness. By leveraging these insights, researchers and practitioners can work towards building more secure and trustworthy machine learning systems that safeguard sensitive information while maintaining robustness to perturbations and adversarial attacks. Future research directions include further exploration of the privacy-robustness trade-off, development of efficient algorithms that achieve both data confidentiality and resilience to perturbations, and investigation of the practical implications of these theoretical findings in real-world machine learning applications. As the field of machine learning continues to evolve, ensuring the privacy and robustness of these algorithms will remain a critical challenge and an area of active research.
Appendix A Relationship between LDP and MBP
In this section, we derive the relationship between LDP and MBP.
A-A LDP implies MBP
The following lemma establishes a connection between Local Differential Privacy (LDP) and Maximum Bayesian Privacy (MBP) under the assumption of a uniform prior belief distribution of the adversary. It states that if a mapping $f_{W|D}(\cdot|\cdot)$ satisfies $\xi$-LDP, then it also guarantees $\xi$-MBP. In other words, if the ratio of the conditional probabilities $f_{W|D}(w|d)$ and $f_{W|D}(w|d^{\prime})$ is bounded within the range $[e^{-\xi},e^{\xi}]$ for any $w\in\mathcal{W}$ and $d,d^{\prime}\in\mathcal{D}$, then the mapping $f_{W|D}(\cdot|\cdot)$ is a privacy-preserving mapping that ensures the log-ratio of the posterior belief $f_{D|W}(d|w)$ and the prior belief $f_{D}(d)$ is bounded by $\xi$ for all $w\in\mathcal{W}$ and $d\in\mathcal{D}$. This lemma provides a useful connection between LDP and MBP, which can be leveraged to design privacy-preserving algorithms.
Lemma A.1 (LDP implies MBP).
Consider the scenario where the prior belief distribution of the adversary is a uniform distribution. That is, $f_{D}(d)=f_{D}(d^{\prime})$ for any $d,d^{\prime}\in\mathcal{D}$. If $f_{W|D}(\cdot|\cdot)$ is $\xi$-LDP. That is, the mapping $f_{W|D}(\cdot|\cdot)$ satisfies that
$\frac{f_{W|D}(w|d)}{f_{W|D}(w|d^{\prime})}\in[e^{-\xi},e^{\xi}],$
then $f_{W|D}(\cdot|\cdot)$ is a privacy-preserving mapping that guarantees $\xi$-maximum Bayesian privacy. That is, $\left|\log\left(\frac{f_{D|W}(d|w)}{f_{D}(d)}\right)\right|\leq\xi$, $\forall w\in\mathcal{W},d\in\mathcal{D}$.
Proof.
From the definition of $\xi$-LDP, we know that
$\frac{f_{W|D}(w|d)}{f_{W|D}(w|d^{\prime})}\in[e^{-\xi},e^{\xi}],$
for all $d,d^{\prime}\in\mathcal{D}$. Notice that
$\frac{f_{W|D}(w|d)}{f_{W|D}(w|d^{\prime})}=\frac{f_{D|W}(d|w)f_{D}(d^{\prime})}{f_{D|W}(d^{\prime}|w)f_{D}(d)}.$
It implies that $\frac{f_{D|W}(d|w)}{f_{D}(d)}\leq\frac{f_{D|W}(d^{\prime}|w)}{f_{D}(d^{\prime})}\cdot e^{\xi}$.
We assume that $f_{D}(d)=f_{D}(d^{\prime})$. Notice that
$\sum_{d\in\mathcal{D}}f_{D|W}(d|w)=1.$
For any fixed $\tilde{d}\in\mathcal{D}$, we have that
11 11 $=\sum_{d\in\mathcal{D}}f_{D|W}(d|w)$
$=f_{D|W}(\tilde{d}|w)+\sum_{d\in\mathcal{D}\setminus\tilde{d}}f_{D|W}(d|w)$
$\leq f_{D|W}(\tilde{d}|w)+e^{\xi}\cdot f_{D|W}(\tilde{d}|w)\cdot(|\mathcal{D}|-1)$
$=(1+e^{\xi}\cdot(|\mathcal{D}|-1))\cdot f_{D|W}(\tilde{d}|w).$
Therefore, we have that
$f_{D|W}(\tilde{d}|w)\geq\frac{1}{1+e^{\xi}\cdot(|\mathcal{D}|-1)}.$
Similarly, we have that
11 11 $=\sum_{d\in\mathcal{D}}f_{D|W}(d|w)$
$=f_{D|W}(\tilde{d}|w)+\sum_{d\in\mathcal{D}\setminus\tilde{d}}f_{D|W}(d|w)$
$\geq f_{D|W}(\tilde{d}|w)+e^{-\xi}\cdot f_{D|W}(\tilde{d}|w)\cdot(|\mathcal{D}|-1)$
$=(1+e^{-\xi}\cdot(|\mathcal{D}|-1))\cdot f_{D|W}(\tilde{d}|w).$
Therefore, we have
$f_{D|W}(\tilde{d}|w)$ $\leq\frac{1}{1+e^{-\xi}\cdot(|\mathcal{D}|-1)}$
$\leq\frac{e^{\xi}}{|\mathcal{D}|}.$
Therefore, $\forall\tilde{d}\in\mathcal{D}$, we have that
$f_{D|W}(\tilde{d}|w)\leq e^{\xi}f_{D}(\tilde{d}).$
We say a mechanism is $\xi$-MBP if $\forall w\in\mathcal{W}^{d}$,
$\frac{f(d|w)}{f(d)}=\frac{f(w|d)f(d)}{f(w)f(d)}=\frac{f(w|d)}{f(w)}\leq e^{\xi}.$ (7)
Therefore, $f_{D|W}(\cdot|\cdot)$ is $\xi$-MBP. ∎
A-B MBP implies LDP
The following lemma illustrates that if an algorithm is $\xi$-maximum Bayesian privacy, then it is $2\xi$-Differentially Private.
Lemma A.2 (MBP implies LDP).
Consider the scenario where the prior belief distribution of the adversary is a uniform distribution. That is, $f_{D}(d)=f_{D}(d^{\prime})$ for any $d,d^{\prime}\in\mathcal{D}$. Let $f_{W|D}(\cdot|\cdot)$ be a privacy-preserving mapping that guarantees $\xi$-maximum Bayesian privacy. That is, $\xi=\max_{w\in\mathcal{W},d\in\mathcal{D}}\left|\log\left(\frac{f_{D|W}(d|w)}{f_{D}(d)}\right)\right|$. Then, the mapping $f_{W|D}(\cdot|\cdot)$ is $2\xi$-Local Differential Private:
$\frac{f_{W|D}(w|d)}{f_{W|D}(w|d^{\prime})}\in[e^{-2\xi},e^{2\xi}].$
Proof.
From the definition of maximum Bayesian privacy leakage, we know that
$\frac{f_{D|W}(d|w)}{f_{D}(d)}\in[e^{-\xi},e^{\xi}],$
for any $w\in\mathcal{W}$ and $d\in\mathcal{D}$. Therefore, for any $d,d^{\prime}\in\mathcal{D}$ we have
$\frac{f_{W|D}(w|d)}{f_{W|D}(w|d^{\prime})}=\frac{f_{D|W}(d|w)f_{D}(d^{\prime})}{f_{D|W}(d^{\prime}|w)f_{D}(d)}\in[e^{-2\xi},e^{2\xi}].$
From the definition of Differential Privacy, we know that it is $2\xi$-Differentially Private. ∎
Combining the results in C-A and C-B, we are now ready to analyze the relationship between MBP and LDP, which is shown in the following theorem.
Theorem A.3 (Relationship between MBP and LDP).
[…]
Appendix B Relationship Between Average Bayesian Privacy and Maximum Bayesian Privacy
Theorem B.1 (Relationship Between MBP and ABP).
Assume that the prior belief of the attacker satisfies that
$\frac{f^{\mathcal{B}}_{D}(d)}{f_{D}(d)}\in[e^{-\epsilon},e^{\epsilon}].$
Let MBP be defined using Definition IV.3. For any $\epsilon_{p,m}\geq 0$, let $f_{W|D}(\cdot|\cdot)$ be a privacy preserving mapping that guarantees $\epsilon_{p,m}$-maximum Bayesian privacy. That is,
$\frac{f_{D|W}(d|w)}{f_{D}(d)}\in[e^{-\epsilon_{p,m}},e^{\epsilon_{p,m}}],$
where $\epsilon_{p,a}$ is introduced in Definition IV.1.
@@ Appendix B Relationship Between Average Bayesian Privacy and Maximum Bayesian Privacy
Combining Eq. (D), Eq. (D) and Eq. (14), we have that

### v5

@@ 4.4 PAC Robustness of Bayesian Private Algorithms
Any $(\epsilon,\beta,\alpha)$-MBP mechanism $\mathcal{M}$ (Definition 4.5) is $\alpha$-PAC-Robust (Definition 4.6) with probability $1-\gamma$ where $\gamma=(1+4\epsilon)\beta$.
This profound result establishes that privacy-preserving mechanisms inherently exhibit robustness to dataset perturbations. The proof (provided in Appendix) shows that the information-limiting properties of private mechanisms constrain how much outputs can vary under input changes, thereby enforcing stability. The robustness confidence parameter $\gamma$ depends explicitly on the privacy parameter $\epsilon$, revealing a quantitative trade-off: stronger privacy ($\epsilon\downarrow$) directly improves robustness guarantees ($\gamma\downarrow$). This provides a principled approach to designing robust algorithms through privacy mechanisms, with applications in adversarial learning and secure data publishing.
The theorem bridges two previously disconnected research domains: privacy preservation and adversarial robustness. By demonstrating that privacy mechanisms inherently provide robustness certificates, it enables joint optimization of both properties. This has significant implications for designing next-generation trustworthy AI systems that require both data protection and operational reliability. The explicit parameter relationships guide practitioners in balancing privacy-robustness trade-offs for specific application requirements.
5 Theoretical Relationship between Data Privacy and Input Robustness
In this section, we introduce the measurements for data privacy and input robustness. These metrics offer a nuanced perspective on the algorithm’s performance in safeguarding sensitive information and its ability to withstand certain distortions.
5.1 Measurements of Data Privacy
First, we introduce the measurement for data privacy.
Definition 5.1 (Data Privacy).
Let $x_{t}^{(i)}$ represent the $i$-th data recovered by the attacker at the $t$-th round of the optimization algorithm, and $x_{o}^{(i)}$ represent the $i$-th original data. The privacy leakage is measured using
$\epsilon_{p}$ $=1-\mathbb{E}\left[\frac{1}{|\mathcal{D}|}\sum_{i=1}^{|\mathcal{D}|}\frac{1}{T}\sum_{t=1}^{T}\frac{||x_{t}^{(i)}-x_{o}^{(i)}||}{D}\right],$ (5)
where the expectation is taken with respect to the randomness of the mini-batch $\mathcal{D}$.
5.2 Measurements of Input Robustness
In assessing the robustness of a machine learning algorithm, we adhere to the following definition.
Definition 5.2 (Input Robustness).
[37, 56] Let $P$ represent a distribution. A model $\mathcal{A}$ is $(r,\epsilon,P)$-input-robust, if it holds that
$\mathbb{E}_{x\sim P}[\text{sup}_{\|\delta\|_{2}\leq r}|\mathcal{A}\left(w,(x+\delta,y)\right)-\mathcal{A}\left(w,(x,y)\right)|]\leq\epsilon,$ (6)
where $x$ denotes the data (training sample or feature), and $y$ denotes the corresponding label.
Remark: The robustness measures the model performance in terms of the data distortion $\|\delta\|$.
5.3 Technical Assumptions
The theoretical analysis relies on three fundamental regularity conditions that establish the mathematical foundation for our privacy-robustness framework. These assumptions are standard in optimization theory yet critical for establishing our main theoretical results.
Assumption 5.1 (L-Lipschitz Gradient Continuity).
The gradient mapping of the loss function $\nabla\mathcal{L}:\mathbb{R}^{d}\to\mathbb{R}^{m}$ satisfies:
$\|\nabla\mathcal{L}(\theta,x)-\nabla\mathcal{L}(\theta,y)\|_{2}\leq L\|x-y\|_{2}$ (7)
for all $x,y\in\mathcal{X}\subset\mathbb{R}^{d}$ and fixed model parameters $\theta$, where $L>0$ denotes the Lipschitz constant.
Remark.
This standard condition in optimization theory ensures bounded gradient variation under input perturbations. The Lipschitz constant $L$ quantifies the maximum rate of gradient change relative to input variations, which is fundamental for convergence analysis and stability guarantees in adversarial settings.
Assumption 5.2 (Bi-Lipschitz Gradient Condition).
There exist constants $0<c_{a}\leq c_{b}<\infty$ such that for all $(x_{1},y),(x_{2},y)\in\mathcal{X}\times\mathcal{Y}$ with $\|x_{1}-x_{2}\|_{2}\leq D$:
$c_{a}\|x_{1}-x_{2}\|_{2}\leq\|\nabla\mathcal{L}(\theta,(x_{1},y))-\nabla\mathcal{L}(\theta,(x_{2},y))\|_{2}\leq c_{b}\|x_{1}-x_{2}\|_{2}$ (8)
where $D$ denotes the maximum pairwise data distance.
Remark.
This condition strengthens Assumption 5.1 by establishing metric equivalence between data space and gradient space [48]. The lower bound ($c_{a}>0$) ensures gradient injectivity necessary for data reconstruction, while the upper bound ($c_{b}$) maintains controlled sensitivity. This holds for strongly convex losses with Lipschitz gradients and prevents pathological landscape scenarios.
Assumption 5.3 (Optimal Regret Bound for Gradient Matching).
The attacker’s reconstruction algorithm achieves $\Theta(\sqrt{T})$ regret:
$c_{0}\sqrt{T}\leq\sum_{t=1}^{T}\|\nabla\mathcal{L}(\theta,(x_{t},y_{t}))-\nabla\mathcal{L}(\theta,(\tilde{x},y_{t}))\|_{2}\leq c_{2}\sqrt{T}$ (9)
where $T$ is the attack horizon, $x_{t}$ the reconstruction at step $t$, and $\tilde{x}$ the target satisfying $\nabla\mathcal{L}(\theta,(\tilde{x},y))=\tilde{g}$.
Remark.
This $\Theta(\sqrt{T})$ regret bound reflects practical attack capabilities with state-of-the-art optimizers:
[…]
where $d$ is data dimension and $\kappa$ the condition number. The regret bound enables tractable analysis while capturing real-world attack efficiency.
5.4 Theoretical Relationship between Data Privacy and Input Robustness
This section establishes a rigorous theoretical connection between data privacy and input robustness in machine learning algorithms. We introduce precise mathematical measures for both privacy leakage and adversarial robustness, providing a unified framework to analyze an algorithm’s dual capabilities in protecting sensitive information while maintaining performance under input perturbations. Our central contribution is a quantitative trade-off theorem demonstrating that privacy-preserving mechanisms fundamentally influence an algorithm’s resilience to adversarial manipulations. This theoretical foundation has profound implications for designing next-generation algorithms that simultaneously optimize both privacy guarantees and robustness properties.
The following lemma establishes a foundational bound on privacy leakage in terms of parameter distortion. It demonstrates that when a semi-honest attacker attempts to reconstruct the original training data from protected model parameters, the achievable privacy leakage is explicitly constrained by the magnitude of intentional distortions introduced during parameter protection.
Lemma 5.4 (The Relationship between Data Privacy and Parameter Distortion [62]).
Assume that the semi-honest attacker employs an optimization algorithm to infer client $k$’s original dataset $d$ from the released parameter $W$. Let $g(d)=W$ represent the parameter generation function, where $\breve{d}$ denotes the dataset corresponding to the original parameter $\breve{W}$, and $g(d)=\partial\mathcal{L}(W,d)/\partial W$ computes the gradient of the loss function. Define the parameter distortion metric as $\Delta=\frac{1}{|\mathcal{D}^{(k)}|}\sum_{m=1}^{|\mathcal{D}^{(k)}|}||g(d^{(m)})-g(\breve{d}^{(m)})||$, where $d^{(m)}$ and $\breve{d}^{(m)}$ denote the $m$-th data points in their respective datasets. When the expected regret of the optimization algorithm over $I>0$ rounds is $\Theta(I^{p})$ and $\Delta\geq\frac{2c_{2}c_{b}}{c_{a}}\cdot I^{p-1}$, the privacy leakage $\epsilon_{p}$ satisfies:
$\epsilon_{p}\leq 1-\frac{c_{a}\cdot\Delta+c_{a}\cdot c_{0}\cdot I^{p-1}}{4D}.$ (10)
This lemma provides a crucial theoretical guarantee: intentional distortion in model parameters directly translates to provable privacy protection. The bound establishes that sufficient parameter distortion ($\Delta$) forces the privacy leakage $\epsilon_{p}$ below a computable threshold, creating a mathematically verifiable shield against training data reconstruction attacks. This formalizes the counterintuitive insight that strategically introduced parameter noise not only preserves privacy but does so in a quantifiable manner governed by the algorithm’s regret behavior and distortion magnitude.
Building upon this foundation, we now present our main theorem which establishes a fundamental connection between privacy leakage and input robustness. The following result characterizes how the privacy-preserving modifications made to an algorithm inherently affect its resilience against input perturbations. Specifically, it derives a functional relationship $f(\epsilon_{p})$ that precisely quantifies how privacy guarantees influence robustness properties.
Theorem 5.5 (Input Robustness of Data Private Algorithms).
Let $\mathcal{A}$ be a protection mechanism with privacy leakage $\epsilon_{p}$ measured according to Definition 5.1, and robustness characterized by Definition 5.2. The robustness parameter $\alpha$ quantifying $\mathcal{A}$’s resilience against adversarial inputs is given by the expression:
$\alpha=\frac{Cr}{2}+\frac{4D(1-\epsilon_{p})-c_{2}c_{b}I^{p-1}}{2c_{a}}$ (11)
where $r$ denotes the robustness radius, $C$ represents the Lipschitz constant, $D$ bounds the parameter space, $\mathcal{D}$ is the dataset domain, and $c_{a}$, $c_{b}$ are algorithm-dependent constants.
This theorem establishes a profound mathematical duality: privacy protection mechanisms inherently enhance algorithmic robustness against input perturbations. The derived expression demonstrates that reduced privacy leakage ($\epsilon_{p}\downarrow$) directly increases the robustness parameter ($\alpha\uparrow$), revealing an inherent synergy between these seemingly competing objectives. This theoretical insight revolutionizes our understanding of privacy-robustness trade-offs, showing that properly designed privacy mechanisms can simultaneously provide both strong privacy guarantees and enhanced adversarial resilience. The precise functional relationship enables principled co-design of algorithms that jointly optimize these critical properties, opening new avenues for developing trustworthy AI systems. (Complete proof and analysis available in G)
6 Conclusions
In this paper, we thoroughly investigated the complex interplay between various privacy notions, namely Local Differential Privacy (LDP), Average Bayesian Privacy (ABP), and Maximum Bayesian Privacy (MBP), and their fundamental connection to algorithmic robustness in machine learning. We integrated established theoretical relationships between LDP and MBP, and further derived novel quantitative bounds connecting MBP and ABP. These insights are crucial for designing effective privacy-preserving algorithms. Furthermore, we forged a strong link between privacy and PAC robust learning. We showed that privacy-preserving algorithms inherently exhibit PAC robustness, and conversely, how principles of robust learning can guide the construction of privacy mechanisms. A key result quantifies how privacy leakage directly impacts input robustness.
Our contributions provide a unified and deeper understanding of the synergies and trade-offs between privacy and robustness. This theoretical framework is vital for building machine learning systems that are not only secure against privacy breaches but also resilient to data perturbations and adversarial attacks. Future research will focus on further exploring the intricate privacy-robustness trade-off landscape, developing more efficient and practical algorithms that simultaneously achieve strong guarantees in both areas, and validating these theoretical findings through extensive empirical studies in real-world, high-stakes applications. As machine learning continues to evolve, ensuring both the privacy and robustness of its algorithms will remain paramount.
Appendix A ACKNOWLEDGMENTS
This work was supported by the National Science and Technology Major Project under Grant 2022ZD0115301.
Appendix B Relationship between LDP and MBP
B.1 Theoretical Relationship between MBP and ABP
In this section, we derive the relationship between MBP and ABP.
Appendix C Relationship Between Average Bayesian Privacy and Maximum Bayesian Privacy
Theorem C.6 (Relationship Between MBP and ABP).
@@ Appendix C Relationship Between Average Bayesian Privacy and Maximum Bayesian Privacy
Let MBP be defined using Definition 4.3. For any $\epsilon_{p,m}\geq 0$, let $f_{W|D}(\cdot|\cdot)$ be a privacy preserving mapping that guarantees $\epsilon_{p,m}$-maximum Bayesian privacy. That is,
@@
where $\epsilon_{p,a}$ is introduced in Definition 4.1.
@@
Combining Eq. (C), Eq. (C) and Eq. (14), we have that